The perimeter of a rectangle is 26 inches. The length is 4 more than twice the width. Find the dimensions of the rectangle.

Answer by lwsshak3(11628)

The perimeter of a rectangle is 26 inches. The length is 4 more than twice the width. Find the dimensions of the rectangle.
***
let x=width
2x+4=length
2(length+width)=perimeter
2(2x+4+x)=26
6x+8=26
6x=18
x=3
2x+4=10
width=3 inches
length=10 inches